# DietGlance: Dietary Monitoring and Personalized Analysis at a Glance with Knowledge-Empowered AI Assistant

### v2

@@ 3 Existing Practices of Dietary Monitoring and Analysis
3.1 Participants and Procedure
We recruited 45 participants with diverse demographic and socio-economic backgrounds by word-of-mouth and online advertisement with inclusion criteria of being adults (aged > 18), fluent in Chinese or English, and without eating disorders. Their ages range from 18 to 62 ($M=33.24,SD=10.62$), including 22 males, 21 females, 1 non-binary, and 1 preferred not to disclose. Participants had diverse socio-economic backgrounds, including twenty university students, six senior researchers, four engineers, one high school graduate, one government employee, one freelancer, one bank employee, two high school teachers, two manufacturers, two workers, one clerk, and four retired. We included both experienced (exp., 23, 51%) and inexperienced (inexp., 22, 49%) participants. The exp. group consisted of individuals who had prior experience using dietary monitoring tools, such as mobile apps, online tools, and paper journals, or receiving professional guidance from experts like dietitians or nutritionists, while the inexp. group included those with no prior experience.
Participants completed a survey on their backgrounds, including experience with dietary monitoring tools, eating environments, and usage frequency, followed by a semi-structured interview on the vision of helpful features, obstacles, and expectations for future eyewear-based dietary systems. The study lasted about 20 minutes, with participants compensated about \$2.5. For qualitative analysis, the lead author first conducted thematic coding and iteratively discussed with other authors to finalize the codes as used in [83].
3.2 Key Findings and Design Requirements
@@
R1. Supporting Automated Accurate Diet Identification and Logging. Participants (30 out of 45) confirmed the need for automated detection and logging of diet consumption to reduce manual effort and improve convenience [5, 70], as the complexity and time demands were the main reasons for both exp. and inexp. participants not or reduce monitoring their diets. Accurate diet identification is crucial for meaningful dietary analysis. Thus, the system should automatically and accurately identify and log both the type and amount of food and drink consumed.
@@
R2. Facilitating Nutritional Analysis based on Reliable Sources. Participants expected professional nutritional information, including detailed nutrient breakdowns, consistent with prior research emphasizing the importance of nutritional analysis [42, 74]. However, they emphasized the inaccurate and incomplete databases in existing tools (9 out of 23 exp.). The system should integrate reliable food databases to support nutrition estimates covering a wide variety of items. Twenty-one out of 45 participants mentioned it would be helpful to track nutrition consumption. This would help users manage their nutrition consumption to maintain health or achieve specific goals.
@@
R3. Providing Personalized Dietary Suggestions with Domain-Specific Knowledge. 33 out of 45 Participants were highly interested in reliable and personalized dietary recommendations tailored to their profiles (e.g., age and weight), goals (e.g., muscle development, weight management, general wellness), and specific dietary needs (e.g., diabetes, allergies, vegetarianism), which is consistent with prior work [74]. Our findings emphasized the need for seamlessly integrating domain knowledge cost-effectively. The system should provide dietary suggestions based on individual dietary behaviors and domain-specific knowledge. Various guidances were desired, including nutrition balance, keeping fit, customized meal plans, ingredient substitutions, portion control, and dietary assessments, etc.
@@
R4. Protecting Privacy. Privacy concerns were raised by 24 participants towards the eyewear-based dietary monitoring, especially regarding the potential of recording their phone screen and conversations with others during meals. Therefore, the system should try to reduce the capture of sensitive information during meal sessions and incorporate the mechanism to protect privacy.
4 System Design and Implementation
@@
Based on the design requirements, we developed DietGlance, as shown in Fig. 2. First, we propose a multimodal sensing approach using camera, IMU, and audio sensors on eyeglasses to detect ingestion behaviors (Sec. 4.1) and capture meal images (Sec. 4.2) (R1, R4). The processed images and user profiles were analyzed using GPT-4V to identify the type and amount of food consumed (Sec. 4.3, R1). Unlike previous studies relying solely on LLM or requiring expert input, we further grounded the LLM with a reliable nutrition library through the RAG module to obtain nutritional analysis and personalized dietary suggestions (Sec. 4.3, R2, R3). Finally, we developed a novel user interface, distinct from previous studies, that allows users to interactively review and correct data, add details, explore results, and obtain more information (Sec. 4.4, R1-R4). All prompts are available in Supplement 1.
@@
Fig. 3: An illustration of the multimodal sensing framework for ingestive episode detection and diet image capture.
4.1 Ingestive Episode Detection
@@
Accurate detection of ingestive episodes is the first step of an automated diet analysis system. DietGlance leveraged IMU and audio signals and built a multimodal machine learning model to achieve the goal.
4.1.1 Data Preprocessing and Feature Extraction
@@
The IMU data used included tri-axial accelerometer and gyroscope data, sampled at 800Hz and 1000Hz from two IMUs on each side of the Aria Glasses. The 800Hz IMU data was upsampled to 1000Hz. Audio signals, captured at 48kHz, were down-sampled to 1000Hz, at which rate the speech data is largely unintelligible, to preserve sensitive content [52] (R4).
@@
IMU and audio signals were normalized and segmented into one-second non-overlapping windows. The IMU data were normalized and concatenated into a 12-channel input (two IMUs, each with a tri-axial accelerometer and a tri-axial gyroscope). The audio signals were processed by Short-Time Fourier Transform (STFT) into Melspectrogram. We further extracted statistical features for each window. For IMU, the mean, standard deviation, min, max, kurtosis, skewness, and frequency features were extracted for the two tri-axial acceleration data and two gyroscope data. To obtain the frequency features, we first computed the Fast Fourier Transform (FFT) of the IMU data and then computed the mean and peak of the magnitude of the frequency domain. For audio signals, we computed the mean, standard deviation of its Melspectrogram, Zero Crossing Rate, Spectral Centroid, Spectral Bandwidth, Spectral Rolloff, Chroma STFT, and Root Mean Square Error. The statistical features of IMU and audio data for each time window were concatenated as the statistical feature.
4.1.2 Ingestive Episode Detection
@@
After feature extraction, we employed a multimodal learning model (Fig. 4), building upon the framework introduced in [52] and further integrating statistical features, to classify time windows as ingestive or non-ingestive activities (R1). Specifically, instead of distinguishing between eating and drinking, we categorized both under the broader term “ingestive behaviors.” This decision accounts for practical challenges, such as the overlap in actions like consuming noodle soup (involving both eating and drinking) or swallowing soft foods without chewing. This simplification aligns with the study’s goals and improves usability in real-world applications.
Fig. 4: The multimodal learning model for classifying ingestive episodes and others.
4.2 Diet Image Capture and Segment
@@
4.2.1 IMU-based Pitch Angle Check
Simultaneously with ingestive episode detection, the pitch angle was calculated from tri-axial acceleration data ($a_{x},a_{y},a_{z}$) as $pitch=arctan(a_{x}/\sqrt{a_{y}^{2}+a_{z}^{2}})$. When the pitch angle exceeded a threshold ($pitch_{th}=5$, indicating the user tilting their head forward), egocentric images were captured at 10fps for three seconds after a behavior transition, ensuring the meal was fully captured (Fig. 3). This approach allowed the system to capture a series of images throughout the meal, including dishes added midway. Although users may tilt their heads back while drinking, capturing beverages from this angle in egocentric images is challenging. By focusing on images captured when users tilted their heads downward, beverages and other meal components were recorded more effectively.
4.2.2 Food, Drink, Tableware Segmentation
We further segmented the food, drinks, and tableware within the images. To exclude the food from nearby diners and include the user’s meal more completely in the egocentric images, we set the bottom one-third region as the primary area of interest (detailed in Supplement 2.1). For segmentation, we employed Grounded-Segment-Anything (Grounded-SAM) [64], which enables precise identification based on textual descriptions. This approach allows the segmentation of a wide range of foods and drinks using class prompts and can easily accommodate new food types by adding new prompts. As illustrated in Fig. 5, we first utilized the pre-trained GroundingDINO model to detect objects based on the class prompt, then applied the pre-trained SAM model to convert detections into masks. Additionally, the class prompt (detailed in Supplement 1.1) included not only food and drink categories but also tableware, as these are often the only remaining objects post-meal.
Fig. 5: An illustration of diet image capture and segment.
4.2.3 Irrelevant Information Blurring and Similarity Check
After segmentation, we blurred the images except for the segmented part with Gaussian Blur [31] to reduce privacy concerns and also highlight the area of interest (i.e., the meal of the user). We then screened these images by comparing the Cosine Similarities of the features extracted from segmented parts using ResNet50 [63] to remove redundancy. If the similarity between the segmented parts was larger than 0.75, we removed one of the images. The blurred part of the images would be further cropped to highlight the meal items and enhance privacy protection. In this way, we finally obtained a series of privacy-preserving meal images.
4.3 Dietary Analysis
@@ 4.3.1 Diet Identification
LLMs, trained on diverse datasets, excel at identifying various cuisines and ingredients in images and estimating food quantities through visual cues in a zero-shot manner, offering detailed nutritional information. In this study, we select GPT-4V among all LLMs due to its superior performance in diet recognition and estimation in open-world settings [42] (prompt available in Supplement 1.2). As shown in Fig. 6 (pink part), the meal log images and user profile (including gender, age, height, and weight) were fed into the GPT-4V to obtain the types and amount of each food and drink consumed by the user in this meal session.
@@
Fig. 6: An illustration of RAG-grounded GPT-4V for diet identification, nutritional analysis, and dietary suggestion.
4.3.2 RAG Module Construction
@@
While LLMs excel in diet identification, they may generate seemingly plausible but inaccurate results, lack health guidelines, and miss domain-specific knowledge. To address this, we integrated RAG to enhance the generative capability of LLM by incorporating knowledge from an extensive nutrition library, thereby providing a more reliable, contextually relevant, and factual response [37].
@@
To build the RAG module, we have constructed a comprehensive nutrition library consisting of documents from a diverse collection of reliable resources. The inclusion criteria included (1) the government-approved food nutrient database, (2) official reports and booklets on food consumption and dietary suggestions for the general public and specific groups, issued by the government or the World Health Organization (WHO), (3) webpages from official government websites, containing recipes, scientific articles about food and nutrition, and healthy diet suggestions, (4) articles published by highly trustworthy organizations (e.g., National Institutes of Health), (5) peer-reviewed research articles from prestigious journals with 28 different needs (e.g., weight management, muscle development, and blood glucose stabilization). Detailed information is available in Supplement 2.2.
@@
We then built the RAG module using LangChain framework [34]. The documents in the nutrition library were first split into chunks with LangChain’s RecursiveCharacterTextSplitter, resulting in 16,438 pages, as shown in Fig. 6 (blue part). Then OpenAI’s embedding model text-embedding-ada-002 created embeddings stored with Facebook AI Similarity Search (Faiss) vector database. Based on these embeddings, the most relevant chunks will be searched based on the query to provide context for answering the query, enhancing the relevance and accuracy of the responses by incorporating the contexts. By invoking the retrieval chain created by LangChain, it would not only return the answer based on the prompt and query but also the sources of the answer from the relevant context, making the sources of dietary suggestions more transparent.
4.3.3 Nutritional Analysis and Dietary Suggestion

### v3

@@ 3. Existing Practices of Dietary Monitoring and Analysis
3.1. Participants and Procedure
We recruited 45 participants with diverse demographic and socio-economic backgrounds by word-of-mouth and online advertisement, with inclusion criteria of being adults (aged ¿ 18), fluent in Chinese or English, and without eating disorders. Eligibility regarding eating disorders was determined based on self-identification during the initial online screening process. The participants were geographically located across three regions: 26 in Mainland China, 17 in Hong Kong SAR, and 2 in Japan. Their ages range from 18 to 62 ($M=33.24,SD=10.62$), including 22 males, 21 females, 1 non-binary, and 1 preferred not to disclose. Participants had diverse socio-economic backgrounds, including twenty university students, six senior researchers, four engineers, one high school graduate, one government employee, one freelancer, one bank employee, two high school teachers, two manufacturers, two workers, one clerk, and four retirees. We included both experienced (exp., 23, 51%) and inexperienced (inexp., 22, 49%) participants. The exp. group consisted of individuals who had prior experience using dietary monitoring tools, such as mobile apps, online tools, and paper journals, or receiving professional guidance from experts like dietitians or nutritionists, while the inexp. group included those with no prior experience.
The procedure for the needs-finding study consisted of two main stages.
  1. (1)
    Online Survey: First, participants were provided with an online questionnaire (Appendix A.1). This survey was used to gather quantitative data on their demographic information, backgrounds, prior experience with dietary monitoring tools, and eating environments.
  2. (2)
    Semi-Structured Interview: After completing the survey, each participant attended a one-on-one semi-structured interview session (conducted either face-to-face or remotely, Appendix A.2), which lasted about 20 minutes per participant. This qualitative session was designed to explore their vision for helpful features, the obstacles they’ve encountered in dietary tracking, and their specific expectations and concerns regarding future eyewear-based dietary systems.
After completing the interview, each participant was compensated about \$2.5. For the subsequent qualitative analysis of the interview data, the lead author conducted thematic coding and iteratively discussed with other authors to finalize the codes.
3.2. Key Findings and Design Requirements
@@
R1. Supporting Automated Accurate Diet Identification and Logging. Participants (30 out of 45) confirmed the need for automated detection and logging of diet consumption to reduce manual effort and improve convenience (Bedri et al., 2020; Shin et al., 2022), as the complexity and time demands were the main reasons for both exp. and inexp. participants not or reduce monitoring their diets. Accurate diet identification is crucial for meaningful dietary analysis. Thus, the system should automatically and accurately identify and log both the type and amount of food and drink consumed.
@@
R2. Facilitating Nutritional Analysis based on Reliable Sources. Participants expected professional nutritional information, including detailed nutrient breakdowns, consistent with prior research emphasizing the importance of nutritional analysis (Lo et al., 2024; Szymanski et al., 2024). However, nine out of 23 exp. participants emphasized the inaccurate and incomplete databases in existing tools. The system should integrate reliable food databases to support nutrition estimates covering a wide variety of items. Twenty-one out of 45 participants mentioned it would be helpful to track nutrition consumption. This would help users manage their nutrition consumption to maintain health or achieve specific goals.
@@
R3. Providing Personalized Dietary Suggestions with Domain-Specific Knowledge. 33 out of 45 participants were highly interested in reliable and personalized dietary recommendations tailored to their profiles (e.g., age and weight), goals (e.g., muscle development, weight management, general wellness), and specific dietary needs (e.g., diabetes, allergies, vegetarianism), which is consistent with prior work (Szymanski et al., 2024). Our findings emphasized the need for seamlessly integrating domain knowledge cost-effectively. The system should provide dietary suggestions based on individual dietary behaviors and domain-specific knowledge. Participants desired guidance on various topics, including nutritional balance, weight management, customized meal plans, ingredient substitutions, portion control, and dietary assessments.
@@
R4. Protecting Privacy. Privacy concerns were raised by 24 participants towards the eyewear-based dietary monitoring, especially regarding the potential of recording their phone screen and conversations with others during meals. Therefore, the system should try to reduce the capture of sensitive information during meal sessions and incorporate a mechanism to protect privacy.
4. System Design and Implementation
@@
Based on the design requirements, we developed DietGlance, as shown in Fig. 2. First, we propose a multimodal sensing approach using camera, IMU, and audio sensors on eyeglasses to detect ingestion behaviors (Sec. 4.1) and capture meal images (Sec. 4.2) (R1, R4). The processed images and user profiles were analyzed using GPT-4V to identify the type and amount of food consumed (Sec. 4.3, R1). Unlike previous studies relying solely on LLM or requiring expert input, we further grounded the LLM with a reliable nutrition library through the RAG module to obtain nutritional analysis and personalized dietary suggestions (Sec. 4.3, R2, R3). Finally, we developed a novel user interface, distinct from previous studies, that allows users to interactively review and correct data, add details, explore results, and obtain more information (Sec. 4.4, R1-R4). All prompts are available in Appendix B.
@@
Fig. 3. An illustration of the multimodal sensing framework for ingestive episode detection and diet image capture. Image capture is triggered for three seconds after both an IMU-based pitch angle check and a behavior transition (e.g., from non-ingestive ’0’ to ingestive ’1’, and vice-versa).
4.1. Ingestive Episode Detection
@@
Accurate detection of ingestive episodes is the first step of an automated diet analysis system. DietGlance leveraged IMU and audio signals and built a multimodal machine learning model to achieve the goal (Fig. 3).
4.1.1. Data Preprocessing and Feature Extraction
@@
The IMU data used included tri-axial accelerometer and gyroscope data, sampled at 800Hz and 1000Hz from two IMUs on each side of the Aria Glasses. The 800Hz IMU data was upsampled to 1000Hz. Audio signals, captured at 48kHz, were down-sampled to 1000Hz, at which rate the speech data is largely unintelligible, to preserve sensitive content (Mollyn et al., 2022) (R4).
@@
IMU and audio signals were normalized and segmented into one-second non-overlapping windows. The IMU data were normalized and concatenated into a 12-channel input (two IMUs, each with a tri-axial accelerometer and a tri-axial gyroscope). The audio signals were processed by Short-Time Fourier Transform (STFT) into Mel-spectrograms. We further extracted statistical features for each window. For IMU, the mean, standard deviation, min, max, kurtosis, skewness, and frequency features were extracted for the two tri-axial acceleration data and two gyroscope data. To obtain the frequency features, we first computed the Fast Fourier Transform (FFT) of the IMU data and then computed the mean and peak of the magnitude of the frequency domain. For audio signals, we computed the mean, standard deviation of their Mel-spectrograms, Zero Crossing Rate, Spectral Centroid, Spectral Bandwidth, Spectral Rolloff, Chroma STFT, and Root Mean Square Error. The statistical features of IMU and audio data for each time window were concatenated as the statistical feature.
4.1.2. Ingestive Episode Detection
As illustrated in Fig. 4, the model consists of three parallel input branches that are processed in two main stages. First, the Audio Feature (Melspectrogram) is processed by a 2D Convolutional Network (a stack of [Conv2d, ReLU, MaxPool2d] repeated 4 times), and the IMU Feature (Acceleration, Gyroscope) is processed by a 1D Convolutional Network ([Conv1d, ReLU, BatchNorm1d, MaxPool2d] repeated 4 times). The outputs of these two convolutional branches are then concatenated and fed into a shared MLP ([Linear, ReLU] repeated 3 times). In parallel, the Statistical Feature is processed by its own MLP ([Linear, ReLU] repeated 4 times). Finally, the outputs of the Audio and IMU MLP branch and the Statistical MLP branch are concatenated and fed into a final MLP ([Linear, ReLU] repeated 2 times). The output layer uses a Sigmoid activation function to produce a probability score between 0 and 1, classifying the window as either non-ingestive (0) or ingestive (1).
@@ 4.1.2. Ingestive Episode Detection
Specifically, instead of distinguishing between eating and drinking, we categorized both under the broader term “ingestive behaviors.” This decision accounts for practical challenges, such as the overlap in actions like consuming noodle soup (involving both eating and drinking) or swallowing soft foods without chewing. This simplification aligns with the study’s goals and improves usability in real-world applications.
Fig. 4. The multimodal learning model for classifying ingestive episodes and others.
4.2. Diet Image Capture and Segment
@@
4.2.1. IMU-based Pitch Angle Check
Simultaneously with ingestive episode detection, we employ a pitch angle check as a practical filter to ensure the egocentric camera is physically pointed at the meal during image capture. The pitch angle was calculated from tri-axial acceleration data ($a_{x},a_{y},a_{z}$) as $pitch=arctan(a_{x}/\sqrt{a_{y}^{2}+a_{z}^{2}})$, where a value exceeding a threshold ($pitch_{th}=5$) indicated a state of being tilted forward. The pitch angle check helps increase the comprehensiveness of the meal capture while also avoiding the capture of the surrounding environment, which benefits privacy. Our image capture is then triggered by a dual-condition logic as illustrated in Fig. 3: images are captured at 10fps for three seconds only when (1) a behavior transition is detected (i.e., from non-ingestive to ingestive, or vice-versa), and (2) the user’s head is in this downward-tilted state. This approach allowed the system to capture a series of images throughout the meal, including dishes added midway.
Although users may tilt their heads back while drinking, capturing beverages from this angle in egocentric images is challenging. By focusing on images captured when users tilted their heads downward, beverages and other meal components were recorded more effectively.
4.2.2. Food, Drink, Tableware Segmentation
We further segmented the food, drinks, and tableware within the images. To exclude the food from nearby diners and include the user’s meal more completely in the egocentric images, we set the bottom one-third region as the primary area of interest (detailed in Appendix C.1). For segmentation, we employed Grounded-Segment-Anything (Grounded-SAM) (Ren et al., 2024), which enables precise identification based on textual descriptions. This approach allows the segmentation of a wide range of foods and drinks using class prompts and can easily accommodate new food types by adding new prompts. As illustrated in Fig. 5, we first utilized the pre-trained GroundingDINO model to detect objects based on the class prompt, then applied the pre-trained SAM model to convert detections into masks. Additionally, the class prompt (detailed in Appendix B.1) included not only food and drink categories but also tableware, as these are often the only remaining objects post-meal.
Fig. 5. An illustration of diet image capture and segmentation.
4.2.3. Irrelevant Information Blurring and Similarity Check
After segmentation, we blurred the images except for the segmented part with Gaussian Blur (Jiang et al., 2023a) to reduce privacy concerns and also highlight the area of interest (i.e., the meal of the user). We then screened these images by comparing the Cosine Similarities of the features extracted from segmented parts using ResNet50 (Rajpal et al., 2021) to remove redundancy. If the similarity between the segmented parts was larger than 0.75, we removed one of the images. The blurred part of the images would be further automatically cropped (to the bounding box of the segmented meal region) to highlight the meal items and enhance privacy protection. In this way, we finally obtained a series of privacy-preserving meal images.
4.3. Dietary Analysis
@@ 4.3.1. Diet Identification
LLMs, trained on diverse datasets, excel at identifying various cuisines and ingredients in images and estimating food quantities through visual cues in a zero-shot manner, offering detailed nutritional information. In this study, we select GPT-4V among all LLMs due to its superior performance in diet recognition and estimation in open-world settings (Lo et al., 2024) (prompt available in Appendix B.2). As shown in Fig. 6 (pink part), the meal log images and user profile (including gender, age, height, and weight) were fed into the GPT-4V to obtain the types and amount of each food and drink consumed by the user in this meal session.
@@
Fig. 6. An illustration of RAG-grounded GPT-4V for diet identification, nutritional analysis, and dietary suggestion.
4.3.2. RAG Module Construction
@@
While LLMs excel in diet identification, they may generate seemingly plausible but inaccurate results, lack health guidelines, and miss domain-specific knowledge. To address this, we integrated RAG to enhance the generative capability of LLM by incorporating knowledge from an extensive nutrition library, thereby providing a more reliable, contextually relevant, and factual response (Lewis et al., 2020).
@@
To build the RAG module, we have constructed a comprehensive nutrition library consisting of documents from a diverse collection of reliable resources. The inclusion criteria included (1) the government-approved food nutrient database, (2) official reports and booklets on food consumption and dietary suggestions for the general public and specific groups, issued by the government or the World Health Organization (WHO), (3) online resources from official government websites, containing recipes, scientific articles about food and nutrition, and healthy diet suggestions, (4) articles published by highly trustworthy organizations (e.g., National Institutes of Health), (5) peer-reviewed research articles from prestigious journals with 28 different needs (e.g., weight management, muscle development, and blood glucose stabilization). These inclusion criteria were designed to ensure the corpus was composed of high-quality, expert-vetted sources to maximize accuracy and minimize the risk of conflicting information. Detailed information is available in Appendix C.2.
@@
We then built the RAG module using the LangChain framework (LangChain Documentation, 2024). The documents in the nutrition library were first split into chunks with LangChain’s RecursiveCharacterTextSplitter, resulting in 16,438 pages, as shown in Fig. 6 (blue part). Then OpenAI’s embedding model text-embedding-ada-002 created embeddings stored with Facebook AI Similarity Search (Faiss) vector database. Based on these embeddings, the most relevant chunks will be searched based on the query to provide context for answering the query, enhancing the relevance and accuracy of the responses by incorporating the contexts. By invoking the retrieval chain created by LangChain, it would not only return the answer based on the prompt and query but also the sources of the answer from the relevant context, making the sources of dietary suggestions more transparent.
4.3.3. Nutritional Analysis and Dietary Suggestion